OpenStudy (anonymous):

have a word problem: Sunday 60 Minutes; Monday 30 Minutes; Tuesday 60 Minutes; Wednesday 0 Minutes; Thursday 60 Minutes; Friday 40 Minutes; Saturday 30 Minutes. How many minutes would it take to practice on Wednesday to average 45 minutes a day?

OpenStudy (anonymous):

(60+30+60+x+60+40+30)/7= 45 280+x=45*7 280+x= 315 x =315-280 x =35

OpenStudy (anonymous):

so, 45minutes would it take to practice on Wednesday to average 45 minutes a day

OpenStudy (anonymous):

35